Backlinks to DiaryRun5 in ARGOS Web (Search all webs)

Results from ARGOS web retrieved at 00:07 (GMT)

Commissioning
Commissioning run diaries Run 1st day Run ID Description 20111128 Run # 1 Nov 2011 Installation of dichroic interfaces 20130614 Run Installati...
DayTime20150213
See e mail from Gustavo: The February 2015 ARGOS campaign started yesterday. Matthias, Mathias and Franz from MPE are on site. The first two days of the campaign...
DayTime20150215
See e mail from Gustavo: All, Work has continued on the DX LM1 actuators by M, M and F. Julian Z. arrived yesterday, and he has been working on the alarm system a...
DayTime20150218
See e mail from Michael L.: Dear Friends of ARGOS, Yesterday (2/18/15) and today (2/19/15), ARGOS activities included the following: Installation and test of the ...
DayTime20150219
Pockels Cells Transmission test Sweep between 0.5 and 2.0us with 4 steps with Internal Sources. Then between 0.1 and 2.0 in 11 steps. The system configuration is ...
DayTime20150220
LM1 calibration DX Measuring the LM1 limit switch position as well as the optical and mechanical zero position. Lower limit AbsEnc IncEnc SW setting 1913 51...
DayTime20150221
LgswUnitDx operation 8:00: we start opertations at horizon to fix the CalUnitSystem DiodeSource problem. It seems related to the low temperature in the C ring rac...
DayTime20150222
LgswUnitDx operation 8:30: we start and we toss light on the LGSW. Using hexapod parameters of yesterday we are not in the LGSW but just in the PatCams. Today the...
r7 - 05 Mar 2015 - 11:22 by TommasoMazzoni
DayTime20150223
We close the loop with LGSW FLAO BIN2 TS. LGSW SLOPE_OFFSET: 20150223_223830 LGSW DARK: 20150223_225520 TN AO rec JCL PCL Res. WF nm rms Notes ...
NightTime20150223
19:00: bad weather, we deploy cal unit TipTiltSystemDx calibrations We start testing the TipTiltSystemDx: we set FLAO BIN4, 50 50 dichroic on FW1, WhiteLightSour...
NightTime20150224
Dome closed because of snow... IM bewteen piezo TT mirror and TS modes This IM is needed to properly offset the piezo mirror that produces the TT modulation on t...
NightTime20150225
Daytime * debug of the NaN PBL in the TruthSensing. * Test of the procedure to fit the PSF on PatCams Nighttime DARK LGSW: 20150226_021930 We set the FLAO cub...
NightTime20150226
Daytime activity DARK LGSW: 20150226_225951 PockelsCellsTransmission test To check the proper behaviour after 10 hours of rest. Scan between 0.1 and 2.0us in 7 s...
r3 - 18 Mar 2015 - 14:30 by TommasoMazzoni
NightTime20150227
Autogain test (QC LGSW) We tested the autogain procedure in different environments: no disturb, disturb =1.2", PYR LGSW and QC LGSW. We first optimized the ...
WebStatistics
%IF{ "istopic '.DefaultWebStatistics'" then="$percentINCLUDE{$quot.DefaultWebStatistics$quot}$percent" else="$percentINCLUDE{$quot.DefaultWebStatistics$quot}$perc...
r398 - 19 Jan 2016 - 08:15 by GillesOrban
Number of topics: 16
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback